OpenStudy (aravindg):

The lines joining the origin to the points of intersection of 2x^2+3xy-4x+1=0 and 3x+y=1 is given by __________

OpenStudy (aravindg):

I tried solving the 2 equations first but it doesnt seem to be do the magic.

OpenStudy (aravindg):

@ganeshie8 Any idea?

OpenStudy (austinl):

So, you want the line that goes through the origin (0,0) and the intersection point of the two given lines (x,y)?

OpenStudy (aravindg):

the "lines"

OpenStudy (austinl):

Oh-ho, silly Austin... one is a quadratic! Gimme a minute to ponder this, I will get right back to you!

OpenStudy (aravindg):

Take your time :)

OpenStudy (austinl):

Do you have a graphing calculator?

OpenStudy (anonymous):

ignore first 3 lines

OpenStudy (aravindg):

But I am not allowed to use it in actual exam. So that doesnt help.

OpenStudy (anonymous):

tell if i have done smthng wrong?

OpenStudy (austinl):

You can't use a TI-83/84 on an exam?

OpenStudy (aravindg):

No calculators allowed. Besides I am looking for a solid answer.

OpenStudy (aravindg):

@divu.mkr Can you explain? I am confused.

OpenStudy (anonymous):

just took the equation of y from 1st equation, putted into 2nd ,got the locus of intersection points....then got the slope for d equation by diff..

OpenStudy (aravindg):

What did you do in the 4th step?

OpenStudy (anonymous):

can you specify by line no. from the top?

OpenStudy (aravindg):

4th line.

OpenStudy (anonymous):

ignoring first 3 lines then i just equated the values of y and not ignoring then i took the slope :D

OpenStudy (aravindg):

Still confused. I got the quadratic 7x^2-x+1=0

OpenStudy (anonymous):

if my algebra is a bit wrong then go by concept :P

OpenStudy (aravindg):

Yeah But I dont know what to do after that.

OpenStudy (anonymous):

slope!!!

OpenStudy (aravindg):

How?

OpenStudy (anonymous):

diff it

OpenStudy (aravindg):

It gives the slope of what?

OpenStudy (austinl):

Ok, we have two different equations. One is of a hyperbola: \(2x^2+3xy-4x+1=0 \) The other is of a line: \(3x+y=1\) We can find the points of intersection through a series of steps. 1) Solve the equation of the line for x. \(3x+y=1\) \(3x=-y+1\) \(\displaystyle x=-\frac{y}{3}+\frac{1}{3}\) 2) Substitute that into the equation for the hyperbola. 3) Expand and group like terms and rewrite the equation. 4) Solve the quadratic equation for y to obtain two solutions. 5) We now substitute the values of y already obtained into the equation we have for x.

OpenStudy (aravindg):

^very long process. I am sure there is a better way out.

OpenStudy (anonymous):

thats where i made a mistake....:D PS- get x by solving 7x^2+x-1=0

OpenStudy (aravindg):

hmm....It is a long process. I am thinking of any easier method.

OpenStudy (anonymous):

i think there is.. but i think i have to remind my notes...

OpenStudy (aravindg):

ok

ganeshie8 (ganeshie8):

2x^2+3xy-4x+1=0 ------(1) 3x+y=1 -------(2) they can have two intersecting points, so the equation for pair of lines through origin wud be :- 2x^2+3xy-4x(3x+y)+1(3x+y)^2=0 ---------------(3)

ganeshie8 (ganeshie8):

proof is trivial :- say A(m, n) is one of the intersection point, then A satisfies (1) and (2) << GIVEN. clearly it satisfies (3) also.

ganeshie8 (ganeshie8):

also, (3) satisfies (0, 0) so origin, and the two intersection points of (1) and (2) lie on the two straight lines represented by (3)

OpenStudy (aravindg):

Ah! That makes all sense. So you just homogenized the given conditions isnt it?

ganeshie8 (ganeshie8):

exactly ! i never learned these in my highschool/college, recently i happened to learn to these while working on few problems of @samigupta8 she did many varieties of these problems and knows much better than me... the internals of theorems involved and stuff behind these...

OpenStudy (aravindg):

Ah thanks a lot ..Now I have a clear idea on how to do these.

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

A solution using Mathematica 8 Home Edition for the calculations is attached.

OpenStudy (aravindg):

Thanks :)